log base 3 (10x^2-x-2)=2+2*log base 3X
solve for x
I tried to change the 2*log base 3x into log base 3x^2.
Then I think i should try and get the 2 left on the right side of the equation to be a log base 3 to allow the log base 3 to cancel completely out.
but I am not sure how to change 2 to a log base 3 if that is the correct step. Any help would be great. Thank you.

Answer by stanbon(57228)

log base 3 (10x^2-x-2)=2+2*log base3(X)
solve for x
----------------------
log3(10x^2-x-2) - 2log3(x) = 2
---
log3(10x^2-x-2) - log3(x^2) = 2
---
log3[(10x^2-x-2/x^2] = 2
----
(10x^2-x-2/x^2) = 9
---
10x^2-x-2 = 9x^2
---
x^2-x-2 = 0
Factor:
(x-2)(x+1) = 0
Positive solution:
x = 2
===================
Cheers,
Stan H.
===================